OpenStudy (anonymous):

What percent of 67 is 33? Round to the nearest tenth of a percent.

OpenStudy (hockeychick23):

33/67= ?

OpenStudy (anonymous):

33/67x100=

OpenStudy (anonymous):

33/67 = 0.49 .49 *100 = 49 49%